OpenStudy (phoenixfire):

How do you row reduce this? No matter what I try I get ugly numbers. \[ \left[\begin{matrix} 22 & -16 & -2 \\ -16 & 12 & 0 \end{matrix}\right] \] The answer is \[ \left[\begin{matrix} 1 & 0 & -3 \\ 0 & 1 & -4 \end{matrix}\right] \]

OpenStudy (anonymous):

\[\begin{bmatrix}22&-16&-2\\-16&12&0\end{bmatrix}\] I'll use \(R_i\) to denote the operations I'd use on row \(i\), and \(R_i\to R_j\) to denote which row I'll replace after each operation. First thing I would do would be to divide everything by the gcf, in effect doing \(\dfrac{1}{2}R_1\to R_1\) and \(\dfrac{1}{4}R_2\to R_2\): \[\begin{bmatrix}11&-8&-1\\-4&3&0\end{bmatrix}\] Following, I'd have \(\dfrac{11}{4}R_2+R_1\to R_2\): \[\begin{bmatrix}11&-8&-1\\-11&\frac{33}{4}&0\end{bmatrix}\] \(R_1+R_2\to R_2\): \[\begin{bmatrix}11&-8&-1\\0&\frac{1}{4}&-1\end{bmatrix}\] \(4R_2\to R_2\): \[\begin{bmatrix}11&-8&-1\\0&1&-4\end{bmatrix}\] \(8R_2+R_1\to R_1\): \[\begin{bmatrix}11&0&-33\\0&1&-4\end{bmatrix}\] And finally, \(\dfrac{1}{11}R_1\to R_1\): \[\begin{bmatrix}1&0&-3\\0&1&-4\end{bmatrix}\]

OpenStudy (phoenixfire):

@SithsAndGiggles Oh wow thank you! I had a very similar result about half way through, i might have messed up a negative sign somewhere though.